OpenStudy (anonymous):

NEARLY FINISHED MY HOMEWORK! WILL GIVE MEDAL! Find k given that 2, k and k^2 - 18 are consecutive geometric terms. Please show working out.

OpenStudy (calculusfunctions):

The nth term of a geometric progression is\[t _{n}=ar ^{n -1}\]where a is the first term, r is the common ratio, and n is the # of terms or the term number.\[r =\frac{ t _{2} }{ t _{1} }=\frac{ t _{3} }{ t _{2} }=\frac{ t _{4} }{ t _{3} }...\]